Two speakers, one directly behind the other, are each generating a 329-Hz sound wave. What is the smallest separation distance between the speakers that will produce destructive interference at a listener standing in front of them? Take the speed of sound to be 344 m/s.
Answer #1

Solution)

Given, Frequency f= 329 Hz

Speed of sound v = 344 m / s

Wavelength λ = v / f = 1.04 m

Now, the smallest separation distance between the speakers which produces destructive interference at a listener standing in front ofthem = λ / 2

= 0.522 m (Ans)

=======